# View Transformation Robustness for Multi-View 3D Object Reconstruction With Reconstruction Error-Guided View Selection

### v1

@@ Abstract
View transformation robustness (VTR) is critical for deep-learning-based multi-view 3D object reconstruction models, which indicates the methods’ stability under inputs with various view transformations. However, existing research seldom focuses on view transformation robustness in multi-view 3D object reconstruction. One direct way to improve the models’ VTR is to produce data with more view transformations and add them to model training. Recent progress on large vision models, particularly Stable Diffusion models, has provided great potential for generating 3D models or synthesizing novel view images with only a single image input. To fully utilize the power of Stable Diffusion models without causing extra inference computation burdens, we propose to generate novel views with Stable Diffusion models for better view transformation robustness. Instead of synthesizing random views, we propose a reconstruction error-guided view selection method, which considers the reconstruction errors’ spatial distribution of the 3D predictions and chooses the views that could cover the reconstruction errors as much as possible. The methods are trained and tested on sets with large view transformations to validate the 3D reconstruction models’ robustness to view transformations. Extensive experiments demonstrate that the proposed method can outperform state-of-the-art 3D reconstruction methods and other view transformation robustness comparison methods. Codel will be available at: https://github.com/zqyq/VTR.
@@ 3.1 Multi-view 3D reconstruction model
A multi-view 3D object reconstruction model $\mathit{f}$ usually consists of the following parts [23]: First, an encoder to extract the feature representation for reconstruction from the image set ($I_{1}$$\dots$,$I_{n}$) of Object $\mathit{I}$; Second, multiple 2D image features or 3D volume features are fed into a merger to fuse features from different views; Third, a decoder to predict the corresponding voxel-base 3D shape $\mathit{V}$ from the features map. The overall prediction process is formulated as: $V=f(I_{1},\cdots,I_{n})$. We adopt two typical kinds of multi-view 3D object reconstruction models, CNNs-based Pix2Vox++ [16] and transformer-based LRGT [17]. Denote the 3D voxel ground truth as $V^{gt}$, and the reconstruction loss is $Loss(V,V^{gt})$ (details in Supp.). Following [16] and [17], we use binary cross entropy loss in Pix2Vox++ while Dice loss [43] in LRGT. The model is first trained on the training set of ‘Aligned’ data.
@@ 3.2 View selection module
The view selection module first projects the 3D reconstruction error $\mathit{E}=\left|V-V^{gt}\right|$ to a series of viewpoints, and then selects viewpoints containing the most reconstruction errors to be fed into the view synthesis module (see Figure. 3). Specifically, We denote $v=[\psi,\theta,\phi]\in\mathbb{R}^{3}$ as the viewpoint parameters, where $[\psi,\theta,\phi]$ is the camera rotation along the z-y-x axes using the Tait-Bryan angles, and roll $\theta$ is set as zero. Then, our target is to estimate the $v=[\psi,\phi]$ from the 3D reconstruction error $\mathit{E}$, which indicates the areas with bad reconstruction results of an object. Naturally, if we could provide more information (views) about the areas with large reconstruction errors and try to rectify these errors, the performance of the reconstruction model could be boosted. Particularly, if a camera view can cover the reconstruction error regions more clearly, it is supposed to aid the robustness of the 3D reconstruction model.
@@ 4.1 Experiment settings
Dataset generation. Since no existing dataset is suitable for evaluating view transformation robustness, we generate a new dataset ShapeNet-VTR, based on ShapeNet [18], which consists of larger viewpoint ranges than the original version via rendering new views from the 3D CAD models in ShapeNet. ShapeNet-VTR consists of 13 categories and 39239 objects in total. The objects in each category are divided according to 7::::1::::2 for training, validation, and testing, respectively. ShapeNet-VTR consists of 3 sets: ‘Aligned’, ‘Hemispherical’, and ‘Spherical’, sharing the same object dividing way but differing in the viewpoint range as shown in Figure 5. ‘Aligned’ set is rendered at a fixed $\phi$ of $60^{\circ}$ and every $15^{\circ}$ in $\psi$; In the ‘Hemispherical’ set, the range of views is set as $\phi$in $[0^{\circ},90^{\circ}]$, $\psi$in $[-180^{\circ},180^{\circ}]$, and thus images are randomly rendered in the upper hemisphere. In ‘Spherical’, the range of views is $\phi$in $[-90^{\circ},90^{\circ}]$, $\psi$in $[-180^{\circ},180^{\circ}]$, and thus images are randomly rendered in the spherical space. In each set, each object contains 24 views but from different camera angle ranges. We train the 3D reconstruction models on the training set of ‘Aligned’ at first and use the models’ performance on ‘Hemispherical’ and ‘Spherical’ sets to evaluate the robustness of 3D object reconstruction models to view transformations.
@@
  • Random: Standard data augmentation with random viewpoints, and the view range is set as $\phi$in $[-90^{\circ},90^{\circ}]$, $\psi$in $[-180^{\circ},180^{\circ}]$.

### v2

@@ Abstract
View transformation robustness (VTR) is critical for deep-learning-based multi-view 3D object reconstruction models, which indicates the methods’ stability under inputs with various view transformations. However, existing research seldom focuses on view transformation robustness in multi-view 3D object reconstruction. One direct way to improve the models’ VTR is to produce data with more view transformations and add them to model training. Recent progress on large vision models, particularly Stable Diffusion models, has provided great potential for generating 3D models or synthesizing novel view images with only a single image input. To fully utilize the power of Stable Diffusion models without causing extra inference computation burdens, we propose to generate novel views with Stable Diffusion models for better view transformation robustness. Instead of synthesizing random views, we propose a reconstruction error-guided view selection method, which considers the reconstruction errors’ spatial distribution of the 3D predictions and chooses the views that could cover the reconstruction errors as much as possible. The methods are trained and tested on sets with large view transformations to validate the 3D reconstruction models’ robustness to view transformations. Extensive experiments demonstrate that the proposed method can outperform state-of-the-art 3D reconstruction methods and other view transformation robustness comparison methods. Code is available at: https://github.com/zqyq/VTR.
@@ 3.1 Multi-view 3D reconstruction model
[…]
@@ 3.2 View selection module
The view selection module first projects the 3D reconstruction error $\mathit{E}=\left|V-V^{gt}\right|$ to a series of viewpoints, and then selects viewpoints containing the most reconstruction errors to be fed into the view synthesis module (see Figure. 3). Specifically, We denote $v=[\psi,\theta,\phi]\in\mathbb{R}^{3}$ as the viewpoint parameters, where $\psi,\theta,\phi]$ is the camera rotation along the z-y-x axes using the Tait-Bryan angles, and roll $\theta$ is set as zero. Then, our target is to estimate the $v=[\psi,\phi]$ from the 3D reconstruction error $\mathit{E}$, which indicates the areas with bad reconstruction results of an object. Naturally, if we could provide more information (views) about the areas with large reconstruction errors and try to rectify these errors, the performance of the reconstruction model could be boosted. Particularly, if a camera view can cover the reconstruction error regions more clearly, it is supposed to aid the robustness of the 3D reconstruction model.
@@ 4.1 Experiment settings
Dataset generation. Since no existing dataset is suitable for evaluating view transformation robustness, we generate a new dataset ShapeNet-VTR, based on ShapeNet [18], which consists of larger viewpoint ranges than the original version via rendering new views from the 3D CAD models in ShapeNet. ShapeNet-VTR consists of 13 categories and 39239 objects in total. The objects in each category are divided according to 7::1::2 for training, validation, and testing, respectively. ShapeNet-VTR consists of 3 sets: ‘Aligned’, ‘Hemispherical’, and ‘Spherical’, sharing the same object dividing way but differing in the viewpoint range as shown in Figure 5. ‘Aligned’ set is rendered at a fixed $\phi$ of $60^{\circ}$ and every $15^{\circ}$ in $\psi$; In the ‘Hemispherical’ set, the range of views is set as $\phi$ $\in$ $0^{\circ},90^{\circ}]$, $\psi$ $\in$ $180^{\circ},180^{\circ}]$, and thus images are randomly rendered in the upper hemisphere. In ‘Spherical’, the range of views is $\phi$ $\in$ $90^{\circ},90^{\circ}]$, $\psi$ $\in$ $180^{\circ},180^{\circ}]$, and thus images are randomly rendered in the spherical space. In each set, each object contains 24 views but from different camera angle ranges. We train the 3D reconstruction models on the training set of ‘Aligned’ at first and use the models’ performance on ‘Hemispherical’ and ‘Spherical’ sets to evaluate the robustness of 3D object reconstruction models to view transformations.
@@
  • Random: Standard data augmentation with random viewpoints, and the view range is set as $\phi$ $\in$ $90^{\circ},90^{\circ}]$, $\psi$ $\in$ $180^{\circ},180^{\circ}]$.